OpenStudy (anonymous):

the distance around earth at the equator is about 40074 kilometers. Find the equatorial diameter of earth to the nearest kilometer. A- 12762 km B- 25512 km C- 62884 km D- 6339700 km

OpenStudy (campbell_st):

well use the circumference formula \[C = \pi \times d\] so you know \[40074 = \pi \times d\] just solve for d

OpenStudy (anonymous):

can u help me solve for d please @campbell_st

OpenStudy (campbell_st):

divide both sides of the equation by pi... that's all you need to do

OpenStudy (anonymous):

ok thanks @campbell_st